# MusicHiFi: Fast High-Fidelity Stereo Vocoding

### v3

@@ Abstract
Diffusion-based audio and music generation models commonly perform generation by constructing an image representation of audio (e.g., a mel-spectrogram) and then convert it to audio using a phase reconstruction model or vocoder. Typical vocoders, however, produce monophonic audio at lower resolutions (e.g., 16-24 kHz), which limits their usefulness. We propose MusicHiFi — an efficient high-fidelity stereophonic vocoder. Our method employs a cascade of three generative adversarial networks (GANs) that convert low-resolution mel-spectrograms to audio, upsamples to high-resolution audio via bandwidth extension, and upmixes to stereophonic audio. Compared to past work, we propose 1) a unified GAN-based generator and discriminator architecture and training procedure for each stage of our cascade, 2) a new fast, near downsampling-compatible bandwidth extension module, and 3) a new fast downmix-compatible mono-to-stereo upmixer that ensures the preservation of monophonic content in the output. We evaluate our approach using objective and subjective listening tests and find our approach yields comparable or better audio quality, better spatialization control, and significantly faster inference speed compared to past work. Sound examples are at https://MusicHiFi.github.io/web/.
@@ I Introduction
Recent generation methods offer an exciting new way to create media content. Diffusion models [1], in particular, have shown great promise for fast, high-quality image generation [2, 3] and are rapidly being adopted for audio and music generation [4, 5, 6, 7, 8, 9, 10, 11]. When used for audio [7, 6, 5, 8], diffusion models are commonly used to generate an image representation of audio (e.g. mel-spectrogram) and then a phase reconstruction model or vocoder is used to convert to audio. This two-stage cascaded approach has been shown beneficial [6], but typically generates low-resolution audio (e.g., mono 16-24 kHz) via a diffusion [12] or generative adversarial network (GAN) vocoder [13, 14, 15].
To improve diffusion-based generation quality for audio and music, a high-fidelity vocoder is essential. Current methods to do so largely come in two forms: bandwidth extension (expansion) and mono-to-stereo upmixing. Bandwidth extension (BWE) increases the frequency range or bandwidth of an audio signal and can be achieved using a source-filter model [16], time-domain [17, 18, 19], or spectral-domain [20, 21, 22, 23, 24] neural network. Mono-to-stereo (M2S) upmixing converts a single-channel audio signal into spatialized (left and right) channels. M2S can be achieved via decorrelation [25] or a recent parametric stereo method [26] that describes audio as a single-channel plus parameters that approx. characterizes the stereo image [27].
@@ IV-B Baselines
For vocoding, we compare against BigVGAN [15] and HiFi-GAN [14] all trained on the same data and input features. Our retrained HiFi-GAN model has 14M params and take 52 GFLOPS for one second of audio. We also train a large HiFi-GAN-large model with 1024 input channels with 55M params, while taking 211 GFLOPS for one second of audio. For BWE, we compare against Aero [23], a recent state-of-the-art BWE method that uses a encoder-decoder architecture with LSTM and temporal-based attention layers with 19M params and takes 85 GFLOPS. We also compare a pre-trained checkpoint of AudioSR [24] for BWE (no training code available). For M2S, we compare against a CPU-only open-source decorrelation method [33, 34], denoted as DSP, that divides the signal into transients, noise, and harmonics and decorrelates non-transient content.
@@ IV-E Subjective evaluation
Figure 2: Subjective listening test violins plots. BWE test (left) and M2S test (right). Test conditions include (a) AudioSR, (b) cascaded MusicHiFi-V and Aero and (c) MusicHiFi-V+BWE (d) MusicHiFi-V+BWE with DSP (e) full MusicHiFi.
@@
For our BWE listening test, we created six test examples from the FMA-small dataset, with each sample being 4 seconds long. Test conditions included (a) AudioSR, (b) cascaded MusicHiFi-V and Aero and (c) MusicHiFi-V+BWE as well as the LA and HA. The goal of this task is to understand the perceptual quality of comparable BWE algorithms. For the M2S evaluation, we prepared twelve listening samples, each 3 seconds long, selected from an internal test dataset instead of the FMA dataset as we found a large number of FMA clips were poorly spatialized. Test conditions included (d) cascaded MusicHiFi-V+BWE with DSP and (e) the full MusicHiFi for M2S comparison as well as the LA and HA. All samples for the two tasks are loudness normalized to -23 dBFS before and after being fed into the cascaded methods. For both tests, 22.05 kHz mono signals were used as the low-anchor (LA) and a hidden reference as high-anchor (HA).
@@ IV-F Subjective evaluation results
For M2S evaluation, we find that our (e) MusicHiFi performs best under different M/S panning coefficients and samples produced from the method (d) MusicHiFi-V+BWE with DSP perform similarly to ours when the energy ratio between mid and side channel is the same (0dB). The difference between our approach and the DSP baseline is statistically significant for side/mid rations 6, 12, and 18 via multiple post-hoc paired t-tests with Bonferroni correction [41].

### v4

@@ Abstract
Diffusion-based audio and music generation models commonly perform generation by constructing an image representation of audio (e.g., a mel-spectrogram) and then convert it to waveform using a phase reconstruction model or vocoder. Typical vocoders, however, produce monophonic audio at lower resolutions (e.g., 16-24 kHz), which limits their usefulness. We propose MusicHiFi — an efficient high-fidelity stereophonic vocoder. Our method employs a cascade of three generative adversarial networks (GANs) that convert low-resolution mel-spectrograms to audio, upsamples to high-resolution audio via bandwidth extension, and upmixes to stereophonic audio. Compared to past work, we propose 1) a unified GAN-based generator and discriminator architecture and training procedure for each stage of our cascade, 2) a new fast, near downsampling-compatible bandwidth extension module, and 3) a new fast downmix-compatible mono-to-stereo upmixer that ensures the preservation of monophonic content in the output. We evaluate our approach using objective and subjective listening tests and find our approach yields comparable or better audio quality, better spatialization control, and significantly faster inference speed compared to past work. Sound examples are at https://MusicHiFi.github.io/web/.
@@ I Introduction
Recent generation methods offer an exciting new way to create media content. Diffusion models [1], in particular, have shown great promise for fast, high-quality image generation [2, 3] and are rapidly being adopted for audio and music generation [4, 5, 6, 7, 8, 9, 10, 11]. When used for audio [7, 6, 5, 8], diffusion models are commonly used to generate an image representation of audio (e.g. mel-spectrogram) and then a phase reconstruction model or vocoder is used to convert to waveform. This two-stage cascaded approach has been shown beneficial [6], but typically generates low-resolution audio (e.g., mono 16-24 kHz) via a diffusion [12] or generative adversarial network (GAN) vocoder [13, 14, 15].
To improve diffusion-based generation quality for audio and music, a high-fidelity vocoder is essential. Current methods to do so largely come in two forms: bandwidth extension (expansion) and mono-to-stereo upmixing. Bandwidth extension (BWE) increases the frequency range or bandwidth of an audio signal and can be achieved using a source-filter model [16], time-domain [17, 18, 19], or spectral-domain [20, 21, 22, 23, 24] neural network. Mono-to-stereo (M2S) upmixing converts a single-channel audio signal into spatialized (left and right) channels. M2S can be achieved via decorrelation [25] or a recent parametric stereo method [26] that describes audio as a single-channel plus parameters that approximately characterizes the stereo image [27].
@@ IV-B Baselines
For vocoding, we compare against BigVGAN [15] and HiFi-GAN [14] all trained on the same data and input features. Our retrained HiFi-GAN model has 14M params and take 52 GFLOPS for one second of audio. We also train a large HiFi-GAN-large model with 1024 input channels with 55M params, while taking 211 GFLOPS for one second of audio. For BWE, we compare against Aero [23], a recent state-of-the-art BWE method that uses a encoder-decoder architecture with LSTM and temporal-based attention layers with 19M params and takes 85 GFLOPS. We also compare a pre-trained checkpoint of AudioSR [24] for BWE (no training code available). For M2S, we compare against a CPU-only open-source implementation [^1] https://github.com/s3a-spatialaudio/s3a-decorrelation-toolbox of decorrelation based method [33, 34], denoted as DSP, that divides the signal into transients, noise, and harmonics and decorrelates non-transient content.
@@ IV-E Subjective evaluation
Figure 2: Subjective listening test violins plots. BWE test (left) and M2S test (right). Test conditions include (a) AudioSR, (b) cascaded MusicHiFi-V and Aero and (c) MusicHiFi-V+BWE (d) full MusicHiFi (e) MusicHiFi-V+BWE with DSP.
@@
For our BWE listening test, we created six test examples from the FMA-small dataset, with each sample being 4 seconds long. Test conditions included (a) AudioSR, (b) cascaded MusicHiFi-V and Aero and (c) MusicHiFi-V+BWE as well as the LA and HA. The goal of this task is to understand the perceptual quality of comparable BWE algorithms. For the M2S evaluation, we prepared twelve listening samples, each 3 seconds long, selected from an internal test dataset instead of the FMA dataset as we found a large number of FMA clips were poorly spatialized. Test conditions included (d) the full MusicHiFi and (e) cascaded MusicHiFi-V+BWE with DSP for M2S comparison as well as the LA and HA. All samples for the two tasks are loudness normalized to -23 dBFS before and after being fed into the cascaded methods. For both tests, 22.05 kHz mono signals were used as the low-anchor (LA) and a hidden reference as high-anchor (HA).
@@ IV-F Subjective evaluation results
For M2S evaluation, we find that our (d) MusicHiFi performs best under different M/S panning coefficients and samples produced from the method (e) MusicHiFi-V+BWE with DSP perform similarly to ours when the energy ratio between mid and side channel is the same (0dB). The difference between our approach and the DSP baseline is statistically significant for side/mid rations 6, 12, and 18 via multiple post-hoc paired t-tests with Bonferroni correction [40].
@@ V Conclusion
VI Acknowledgement
Z. Duan would like to thank support from National Science Foundation grant No. 1846184.